OpenStudy (anonymous):

Give the function G(v)= 7v+1,Find each of the following. G(2) G(-6) G(6.3)

OpenStudy (anonymous):

what is 7v+1 when you plug in a 2 for the V?

jhonyy9 (jhonyy9):

G(2)=7*2 +1 =15 G(-6)=7(-6)+1 =-42+1=-41 G(6,3)=7(6,3)+1 =44,1+1 =45,1